A. What is the value today of a money machine that will pay $3,781.00 per year for 12.00 years? Assume the first payment is made 2.00 years from today and the interest rate is 14.00%.

B. What is the value today of a money machine that will pay $4,451.00 every six months for 24.00 years? Assume the first payment is made 5.00 years from today and the interest rate is 7.00%.
Answer #1

A). PV of Annuity = Annual Payment * [{1 - (1 + r)-n} / r]

= $3,781 * [{1 - (1 + 0.14)-12} / 0.14]

= $3,781 * [0.7924 / 0.14]

= $3,781 * 5.6603 = $21,401.56

Value of Money Machine today = PV of Annuity / (1 + r)n

= $21,401.56 / (1 + 0.14) = $18,773.30

B). PV of Annuity = Annual Payment * [{1 - (1 + r)-n} / r]

= $4,451 * [{1 - (1 + 0.07)-24} / 0.07]

= $4,451 * [0.8029 / 0.07]

= $4,451 * 11.4693 = $51,050.01

Value of Money Machine today = PV of Annuity / (1 + r)n

= $51,050.01 / (1 + 0.07)4 = $51,050.01 / 1.3108 = $38,945.80